Wednesday, June 9, 2010

Summary Night Report 20100609

Report for night 2010-06-09

SA: Alexei
SO: Patrick
Others: Sharl, Janus, Anthony

Notes and messages


1. Tracker was tested as the first step
and tracker+structure+dome (but without shutter) after that.

No any major problems were detected.

2. Tracker moves in X,Y directions were tested first and
normal observation mode after that, where objects for pointing were
selected using OPTI, sent to TCS and telescope started after that.

No any major problems were detected.

Small problem with OPTI was detected (for Encarni):
The last version of OPTI takes into account
proper motions in the different way compare to TCS.
For this reason, objects selected with OPTI were not
accepted by TCS finally.
Please, check.

3. 10 pointings to the random objects were done with
calculation of time for each. See log for details.

4. ELS display does not show RA and DEC.
Needs to be fixed.

No comments:

Post a Comment